OpenStudy (dezzierae):

An invasive species of rodent is spreading across the south. The initial population was estimated at 400 individuals in 2010, but the population has been growing by 58% every year. How many of these rodents can we expect by the year 2025 if the rate they are spreading remains the same? 1,871 871 81,871 381,871

OpenStudy (holsteremission):

If \(y_n\) is the population at the \(n\)th year, and the population grows by a factor of \(58\%\) each year, then you can model the population recursively by \[\begin{cases} y_0=400\\[1ex] y_n=1.58y_{n-1}&\text{for }n\ge1 \end{cases}\]where \(n=0\) corresponds to the year \(2010\). You can solve for \(y_n\) inductively: \[y_n=1.58y_{n-1}=1.58^2y_{n-2}=1.58^3y_{n-3}=\cdots=1.58^ny_0\]The population in \(2025\) would be given by \(y_{25}\).